I have a few problems I need help with. Can you help me?All of these are multication and divison
(3x+4)(x^2-5+1)
(1/2x + 1/3)(1/2x-1/4)
(12y^3+18y^2+24y)/(6y)
(2x^3-9x^2+11x-3)/(2x-3)

Answer by stanbon(48545)

(3x+4)(x^2-5+1) = (3x+4)(x^2-4) = 3x^3+4x^2-3x-16
(1/2x + 1/3)(1/2x-1/4) = (1/4)x^2-(1/8)x+(1/6)x-1/12 = (1/4)x^2 +(1/24)x-(1/12)
(12y^3+18y^2+24y)/(6y) = 2y^2+3y+4
(2x^3-9x^2+11x-3)/(2x-3)= x^2-3x+10 with a remainder of 27
===================================
Cheers,
Stan H.